OpenStudy (anonymous):

tan^2x=sec^2-sin^2x-cos^2x

OpenStudy (anonymous):

First use the Pythagorean identity sin^2 (angle) + cos^2 (angle) = 1. (tan^2 2x)+(sin^2 2x)+(cos^2 2x) = (sec^2 2x) (tan^2 2x)+1 = (sec^2 2x) Next another Pythagorean identity is that tan^2 (angle) + 1 = sec^2 (angle), so substitute... sec^2 2x = sec^2 2x Done!